the sum of the age of the mother and daughters is 50 years. also the age of the mother is 7 times the daughter's age 5 years ago.then find their ages now?
Answer by Cromlix(4381)

Sum of mother and daughter's ages = 50 years
M + D = 50 (M = 50 - D)
Age of mother is 7 x daughter's age 5 years ago.
M - 5 = 7(D - 5)
Replace M with 50 - D from first
equation
50 - D - 5 = 7(D - 5)
50 - D - 5 = 7D - 35
-8D = -80
8D = 80
D = 10
Mother = 40 years old
Daughter = 10 years old.
Hope this helps.
:-)